2. (10 points) A cannon fires a cannonball with a final launch speed of 43. m/s. The barrel length of the cannon, over which the cannonball is accelerated, is Determine the acceleration the cannonball experiences while it is being fired. (Assume the cannonball starts from rest.) (E): 192. m/ s (D): 185. m/s (C): 182. m/s (B): 175. m/ s (A): 187. m/s Thanks :)
Answer #1

u =0 , v = 43 m/s, s = 5 m

from kinematic equaitons

v^2 - u^2 = 2as

43^2 -0 = 2*a*5

a = 185 m/s^2

correct option is (D)